# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 3.1 The forcing, group action, and the countable-support filter
Let $\mathscr{G}:=\operatorname{Sym}(\omega_{1})$ be the full permutation group of $\omega_{1}$. Each $\pi\in\mathscr{G}$ acts on $\mathbb{P}$ by
$(\pi p)(\pi(\alpha),n)\;=\;p(\alpha,n),$
equivalently $(\pi p)=\{(\pi(\alpha),n,i):(\alpha,n,i)\in p\}$. This extends (recursively) to an action on $\mathbb{P}$-names in the usual way.
@@
We call $\dot{x}$ symmetric if $\operatorname{Stab}(\dot{x})\in\mathscr{F}$. Let $\operatorname{HS}$ be the class of hereditarily symmetric names: $\dot{x}\in\operatorname{HS}$ iff $\dot{x}$ is symmetric and every name appearing in $\dot{x}$ is in $\operatorname{HS}$.
Fix $G\subseteq\mathbb{P}$ $V$-generic. The associated symmetric extension is
$\mathcal{N}\;:=\;\operatorname{HS}^{G}\;=\;\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}.$
@@ 3.2 CCC and dependent choice
Let $\{p_{\xi}:\xi<\omega_{1}\}\subseteq\operatorname{Add}(\omega,\omega_{1})$ be uncountable. Write $D_{\xi}=\operatorname{dom}(p_{\xi})\subseteq\omega\times\omega_{1}$, so each $D_{\xi}$ is finite.
@@
Let $\langle\mathbb{P},\mathscr{G},\mathscr{F}\rangle$ be a symmetric system, i.e. $\mathscr{G}\leq\operatorname{Aut}(\mathbb{P})$ and $\mathscr{F}$ is a normal filter of subgroups of $\mathscr{G}$. If $G\subseteq\mathbb{P}$ is $V$-generic, then $\operatorname{HS}^{G}=\{\dot{x}^{G}:\dot{x}\in\operatorname{HS}\}$ is a transitive model of $\mathsf{ZF}$ and $V\subseteq\operatorname{HS}^{G}\subseteq V[G]$.
@@
This is the standard symmetric extension theorem; see, e.g., Jech chapter 15, Lemma 15.51[8]. ∎
@@
$\mathsf{ZF}$. By Lemma 3.8, the symmetric extension $\mathcal{N}=\operatorname{HS}^{G}$ is a transitive model of $\mathsf{ZF}$ (containing all ordinals, as usual for symmetric extensions). We record the following symmetry lemma for the action of $\mathscr{G}$ on $\mathbb{P}$ for later use.
@@
$\mathsf{DC}$. Let $A,R\in\mathcal{N}$ with $A\neq\emptyset$ and $\forall x\in A\,\exists y\in A\,(xRy)$. Fix $\dot{A},\dot{R}\in\operatorname{HS}$ with $\dot{A}^{G}=A$ and $\dot{R}^{G}=R$. Choose $p_{\ast}\in\mathbb{P}$ such that
@@ 3.3 Countable supports for $\operatorname{HS}$-names
Notation 3.12 (Restrictions to $E$).
For $E\subseteq\omega_{1}$ and $p\in\operatorname{Add}(\omega,\omega_{1})$, let
$p\!\restriction E\;:=\;p\cap(E\times\omega\times 2),$
be viewed as a condition in $\mathbb{P}_{E}:=\operatorname{Add}(\omega,E)$.
For a $\mathbb{P}$-name $\tau$, define recursively the restricted name $\tau\!\restriction E$ by
$\tau\!\restriction E\;:=\;\{\langle\sigma\!\restriction E,\;p\!\restriction E\rangle:\langle\sigma,p\rangle\in\tau\}.$
Then $\tau\!\restriction E$ is a $\mathbb{P}_{E}$-name (via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$).
Definition 3.13 (Support).
If $\dot{x}\in\operatorname{HS}$, we say that $E\subseteq\omega_{1}$ is a support for $\dot{x}$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
@@
If $\dot{x}\in\operatorname{HS}$, then in particular $\dot{x}$ is $\mathscr{F}$–symmetric, so $\operatorname{Stab}(\dot{x})\in\mathscr{F}$. Since $\operatorname{Stab}(\dot{x})\in\mathscr{F}$, Lemma 3.4 yields a countable $E$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$. This is exactly the assertion that $\dot{x}$ has countable support. ∎
@@
Let $E\subseteq\omega_{1}$ and let $\tau$ be a $\mathbb{P}$-name such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Let $\varphi$ be any formula whose parameters are $\mathbb{P}$-names also fixed by $\operatorname{Fix}(E)$. If $p\Vdash\varphi(\tau)$, then $p\!\restriction E\Vdash\varphi(\tau)$.
@@
Suppose $p\Vdash\varphi(\tau)$ but $p\!\restriction E\nVdash\varphi(\tau)$. Then there is $q\leq p\!\restriction E$ with $q\Vdash\neg\varphi(\tau)$.
Let $F_{p}\subseteq\omega_{1}\setminus E$ be the finite set of ordinals mentioned by $p$ outside $E$ (i.e. those $\alpha\notin E$ with $(\alpha,n)\in\operatorname{dom}(p)$ for some $n$), and define $F_{q}$ similarly. Choose a permutation $\pi\in\operatorname{Fix}(E)$ such that $\pi[F_{p}]$ is disjoint from $F_{q}$ and such that $\pi$ fixes every element of $F_{q}$. Then $\pi p$ is compatible with $q$ (they agree on $E$ and have disjoint requirements outside $E$), so let $r\leq\pi p,q$.
By the symmetry lemma (Claim 3.2) and $\pi\in\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$, we have $\pi p\Vdash\varphi(\pi\tau)=\varphi(\tau)$, hence $r\Vdash\varphi(\tau)$. But also $r\leq q$ and $q\Vdash\neg\varphi(\tau)$, so $r\Vdash\neg\varphi(\tau)$, a contradiction. ∎
@@
The condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ ($\tau$ has symmetry support EE”) means that $\tau$ is invariant under permutations fixing $E$ pointwise. This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G}$ lies in $V[G_{E}]$.
Indeed, letting $\dot{A}=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$. Nevertheless, $\dot{A}$ is not a $\mathbb{P}_{E}$–name for any countable $E$, and $A\notin V[G_{E}]$ for every countable $E\subseteq\omega_{1}$.
@@
Let $E\subseteq\omega_{1}$, let $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$, and let $\tau$ be a $\mathbb{P}_{E}$–name (viewed as a $\mathbb{P}$–name via the inclusion). If $G\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G\cap\mathbb{P}_{E}$, then
$\tau^{G}=\tau^{G_{E}}\in V[G_{E}].$
@@
By induction on the rank of $\tau$. Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G}$ only consults conditions from $G$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$. Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G}=\tau^{G_{E}}$. ∎
@@ 3.4 The canonical Cohen set $A$ and failure of choice
For each $\alpha<\omega_{1}$, define the canonical $\mathbb{P}$-name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real by
$\dot{c}_{\alpha}\;:=\;\{(\check{n},p):p(\alpha,n)=1\}.$
Let $c_{\alpha}:=\dot{c}_{\alpha}^{G}\in 2^{\omega}$.
@@
$A\;:=\;\{c_{\alpha}:\alpha<\omega_{1}\}.$
@@
For each $\alpha<\omega_{1}$, the canonical name $\dot{c}_{\alpha}$ for the $\alpha$-th Cohen real has support $\{\alpha\}$, hence is hereditarily symmetric. Let
$\dot{A}=\{(\dot{c}_{\alpha},\mathbbm{1}_{\mathbb{P}}):\alpha<\omega_{1}\}.$
Any $\pi\in\mathscr{G}$ permutes the $\dot{c}_{\alpha}$’s, so $\pi(\dot{A})=\dot{A}$. Thus $\operatorname{Stab}(\dot{A})=\mathscr{G}\in\mathscr{F}$, and $\dot{A}$ is symmetric. Since its members are hereditarily symmetric, $\dot{A}\in\operatorname{HS}$, hence $A=\operatorname{val}(\dot{A},G)\in\mathcal{N}$. Indeed $\dot{c}_{\alpha}$ is a $\mathbb{P}_{\{\alpha\}}$-name, hence $\dot{c}_{\alpha}\in\operatorname{HS}$ by Lemma 3.15. ∎
@@
Suppose toward a contradiction that there are an ordinal $\theta$ and $\dot{f}\in\operatorname{HS}$ such that $\dot{f}^{G}$ is an injection $A\to\theta$ in $\mathcal{N}$. Choose $p_{0}\in G$ and $\theta$ with
@@
Remark 3.22 (Dedekind-infinitude).
Unlike the finite-support generalized Cohen model (which produces a Dedekind-finite $A$; cf. [15, §2.4]), the present countable-support filter allows many countable subsets of $A$: for each countable $E\subseteq\omega_{1}$ in $V$, the name $\dot{A}_{E}:=\{(\dot{c}_{\alpha},\mathbf{1}):\alpha\in E\}$ is hereditarily symmetric, so $A_{E}:=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}$.
@@
Lemma 3.24 (A forced nonempty name yields a ground-coded witness).
Let $\dot{X}$ be a $\mathbb{P}$–name and suppose $p\Vdash``\dot{X}\neq\varnothing$". Then there exist $q\leq p$ and a $\mathbb{P}$–name $\dot{x}$ such that $q\Vdash``\dot{x}\in\dot{X}$".
Proof.
This is the standard forcing-theorem witness extraction for $\exists$. From $p\Vdash\exists x\,(x\in\dot{X})$ we obtain $q\leq p$ and a name $\dot{x}$ with $q\Vdash\dot{x}\in\dot{X}$. ∎
@@
Suppose $\sigma$ is an order automorphism of $(X,<)$ and $\sigma\neq\mathrm{id}_{X}$. Let $x_{0}$ be the <<-least element of $X$ with $\sigma(x_{0})\neq x_{0}$.
If $\sigma(x_{0})<x_{0}$, then $\sigma(x_{0})<x_{0}$ implies (by minimality of $x_{0}$) that $\sigma(\sigma(x_{0}))=\sigma(x_{0})$. But order preservation yields $\sigma(\sigma(x_{0}))<\sigma(x_{0})$, a contradiction.
If $\sigma(x_{0})>x_{0}$, then apply the previous case to $\sigma^{-1}$: since $\sigma^{-1}$ is also an order automorphism and $\sigma^{-1}(x_{0})<x_{0}$ (because $x_{0}<\sigma(x_{0})$), we get a contradiction.
Hence $\sigma=\mathrm{id}_{X}$. ∎

### v4

@@ 3.1 The forcing, group action, and the countable-support filter
Let $\mathscr{G}:=\operatorname{Sym}(\omega_{1})$ be the full permutation group of $\omega_{1}$. Each $\pi\in\mathscr{G}$ induces an automorphism of $\mathbb{P}$ by permuting the $\omega_{1}$-coordinate:
$\operatorname{dom}(\pi p)=\{(\pi(\alpha),n):(\alpha,n)\in\operatorname{dom}(p)\}\quad\text{and}\quad(\pi p)(\pi(\alpha),n)=p(\alpha,n).$
This extends to $\mathbb{P}$-names by the standard recursion:
$\pi\dot{x}\;:=\;\{\langle\pi\dot{y},\pi p\rangle:\langle\dot{y},p\rangle\in\dot{x}\}.$
@@
For $E\subseteq\omega_{1}$, we say that $E$ supports $\dot{x}$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$.
We call $\dot{x}$ symmetric if $\operatorname{Stab}(\dot{x})\in\mathscr{F}$, equivalently if $\dot{x}$ has some $V$-countable support $E\in[\omega_{1}]^{\leq\omega}$ (by Lemma 3.4).
Let $\operatorname{HS}$ be the class of hereditarily symmetric names: $\dot{x}\in\operatorname{HS}\iff\dot{x}$ is symmetric and every name appearing in $\dot{x}$ lies in $\operatorname{HS}$.
Fix $G_{0}\subseteq\mathbb{P}$ $V$-generic. The associated symmetric extension is
$\mathcal{N}\;:=\;\operatorname{HS}^{G_{0}}\;=\;\{\dot{x}^{G_{0}}:\dot{x}\in\operatorname{HS}\}.$
@@ 3.2 CCC and dependent choice
Let $\{p_{\xi}:\xi<\omega_{1}\}\subseteq\operatorname{Add}(\omega,\omega_{1})$ be uncountable. Write $D_{\xi}=\operatorname{dom}(p_{\xi})\subseteq\omega_{1}\times\omega$, so each $D_{\xi}$ is finite.
@@
Let $\langle\mathbb{P},\mathscr{G},\mathscr{F}\rangle$ be a symmetric system, i.e. $\mathscr{G}\leq\operatorname{Aut}(\mathbb{P})$ and $\mathscr{F}$ is a normal filter of subgroups of $\mathscr{G}$. If $H\subseteq\mathbb{P}$ is $V$-generic, then $\operatorname{HS}^{H}=\{\dot{x}^{H}:\dot{x}\in\operatorname{HS}\}$ is a transitive model of $\mathsf{ZF}$ and $V\subseteq\operatorname{HS}^{H}\subseteq V[H]$.
@@
This is the standard symmetric extension theorem; see [8, Lemma 15.51]. ∎
@@
$\mathsf{ZF}$. By Lemma 3.8, the symmetric extension $\mathcal{N}=\operatorname{HS}^{G_{0}}$ is a transitive model of $\mathsf{ZF}$ (containing all ordinals, as usual for symmetric extensions). We record the following symmetry lemma for the action of $\mathscr{G}$ on $\mathbb{P}$ for later use.
@@
$\mathsf{DC}$. Let $A,R\in\mathcal{N}$ with $A\neq\emptyset$ and $\forall x\in A\,\exists y\in A\,(xRy)$. Fix $\dot{A},\dot{R}\in\operatorname{HS}$ with $\dot{A}^{G_{0}}=A$ and $\dot{R}^{G_{0}}=R$. Choose $p_{\ast}\in\mathbb{P}$ such that
@@ 3.3 Countable supports for $\operatorname{HS}$-names
Notation 3.12 (Coordinate restriction to $E$).
Let $E\subseteq\omega_{1}$. Set
$\mathbb{P}_{E}\;:=\;\operatorname{Add}(\omega,E)\;=\;\operatorname{Fn}(E\times\omega,2,{<}\omega),$
viewed as a (complete) subposet of $\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$ via the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$.
For $p\in\mathbb{P}$, define the restriction of $p$ to $E$ by
$p\!\restriction E\;:=\;p\cap\bigl((E\times\omega)\times 2\bigr).$
Equivalently, $\operatorname{dom}(p\!\restriction E)=\operatorname{dom}(p)\cap(E\times\omega)$ and $(p\!\restriction E)(\alpha,n)=p(\alpha,n)$ for $(\alpha,n)\in\operatorname{dom}(p)\cap(E\times\omega)$. Then $p\!\restriction E\in\mathbb{P}_{E}$ and $p\leq p\!\restriction E$ (reverse inclusion order).
Remark 3.13 (Symmetry supports).
Recall that $E\subseteq\omega_{1}$ supports a $\mathbb{P}$-name $\tau$ if $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. We refer to such $E$ as a symmetry support to distinguish it from forcing supports.
@@
If $\dot{x}\in\operatorname{HS}$, then in particular $\dot{x}$ is $\mathscr{F}$–symmetric, so $\operatorname{Stab}(\dot{x})\in\mathscr{F}$. By Lemma 3.4, there exists a countable $E\in[\omega_{1}]^{\leq\omega}$ with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{x})$. This is exactly the assertion that $\dot{x}$ has countable support. ∎
@@
Lemma 3.16, Remark 3.17, and Lemma 3.18 record standard support and subextension facts for the Cohen seed, and will not be invoked explicitly later; they are included to fix conventions and prevent common confusions about symmetry supports versus $\mathbb{P}_{E}$-names.
@@
Let $E\subseteq\omega_{1}$ and let $\tau$ be a $\mathbb{P}$-name with $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$. Let $\varphi$ be any formula and let $\vec{\sigma}$ be a tuple of $\mathbb{P}$-names such that $\operatorname{Fix}(E)\leq\operatorname{Stab}(\sigma_{i})$ for each $i$. Then for every $p\in\mathbb{P}$,
$p\Vdash\varphi(\tau,\vec{\sigma})\quad\Longleftrightarrow\quad p\!\restriction E\Vdash\varphi(\tau,\vec{\sigma}).$
@@
The implication $\Leftarrow$” holds since $p\leq p\!\restriction E$. For $\Rightarrow$”, suppose $p\Vdash\varphi(\tau,\vec{\sigma})$ but $p\!\restriction E\nVdash\varphi(\tau,\vec{\sigma})$. Then there is $q\leq p\!\restriction E$ with $q\Vdash\neg\varphi(\tau,\vec{\sigma})$.
Let
$\begin{split}&F_{p}:=\{\alpha\in\omega_{1}\setminus E:\exists n\in\omega\ ((\alpha,n)\in\operatorname{dom}(p))\},\\ &F_{q}:=\{\alpha\in\omega_{1}\setminus E:\exists n\in\omega\ ((\alpha,n)\in\operatorname{dom}(q))\}.\end{split}$
Choose $\pi\in\operatorname{Fix}(E)$ such that $\pi[F_{p}]\cap F_{q}=\varnothing$ and $\pi$ fixes every element of $F_{q}$. Then $\pi p$ is compatible with $q$, so pick $r\leq\pi p,q$.
Since $\pi\in\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ and $\operatorname{Fix}(E)\leq\operatorname{Stab}(\sigma_{i})$ for each $i$, we have $\pi\tau=\tau$ and $\pi\sigma_{i}=\sigma_{i}$ for all $i$. By automorphism invariance of the forcing relation, $\pi p\Vdash\varphi(\pi\tau,\pi\vec{\sigma})$, i.e. $\pi p\Vdash\varphi(\tau,\vec{\sigma})$, hence $r\Vdash\varphi(\tau,\vec{\sigma})$. But $r\leq q$ and $q\Vdash\neg\varphi(\tau,\vec{\sigma})$, contradiction. ∎
@@
The condition $\operatorname{Fix}(E)\leq\operatorname{Stab}(\tau)$ ($\tau$ has symmetry support EE”) means that $\tau$ is invariant under permutations fixing $E$ pointwise. This does not imply that $\tau$ is a $\mathbb{P}_{E}$–name, nor that $\tau^{G_{0}}$ lies in $V[G_{E}]$.
Indeed, letting $\dot{A}:=\{\langle\dot{c}_{\alpha},\mathbf{1}\rangle:\alpha<\omega_{1}\}$ be the canonical name for the Cohen set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$, we have $\operatorname{Stab}(\dot{A})=\operatorname{Sym}(\omega_{1})$, so $\dot{A}$ has symmetry support $\emptyset$. Nevertheless, $\dot{A}$ is not a $\mathbb{P}_{E}$–name for any countable $E$, and $A\notin V[G_{E}]$ for every countable $E\subseteq\omega_{1}$.
@@
Let $E\subseteq\omega_{1}$, let $\mathbb{P}_{E}=\operatorname{Add}(\omega,E)\leq\mathbb{P}=\operatorname{Add}(\omega,\omega_{1})$, and let $\tau$ be a $\mathbb{P}_{E}$–name (viewed as a $\mathbb{P}$–name via the inclusion). If $G_{0}\subseteq\mathbb{P}$ is VV–generic and $G_{E}:=G_{0}\cap\mathbb{P}_{E}$, then
$\tau^{G_{0}}=\tau^{G_{E}}\in V[G_{E}].$
@@
By induction on the rank of $\tau$. Using the inclusion $\mathbb{P}_{E}\hookrightarrow\mathbb{P}$, the valuation recursion for $\tau^{G_{0}}$ only consults conditions from $G_{0}$ that appear in $\tau$, and these conditions all lie in $\mathbb{P}_{E}$, hence belong to $G_{E}$. Applying the inductive hypothesis to subnames of $\tau$ yields $\tau^{G_{0}}=\tau^{G_{E}}$. ∎
@@ 3.4 The canonical Cohen set $A$ and failure of choice
Recall that for each $\alpha<\omega_{1}$, $\dot{c}_{\alpha}$ denotes the canonical $\mathbb{P}$-name for the $\alpha$-th Cohen real and $c_{\alpha}:=\dot{c}_{\alpha}^{G_{0}}\in 2^{\omega}$.
@@
Define the canonical $\mathbb{P}$-name
$\dot{A}\;:=\;\{\langle\dot{c}_{\alpha},\mathbf{1}_{\mathbb{P}}\rangle:\alpha<\omega_{1}\},$
and set
$A\;:=\;\dot{A}^{G_{0}}\;=\;\{c_{\alpha}:\alpha<\omega_{1}\}\subseteq 2^{\omega}.$
@@
By Lemma 3.15, each $\dot{c}_{\alpha}$ lies in $\operatorname{HS}$ (indeed it has symmetry support $\{\alpha\}$). For any $\pi\in\mathscr{G}$, $\pi\dot{A}=\dot{A}$, hence $\operatorname{Stab}(\dot{A})=\mathscr{G}\in\mathscr{F}$ and $\dot{A}$ is symmetric. Since all members of $\dot{A}$ are hereditarily symmetric, $\dot{A}\in\operatorname{HS}$. Therefore $A=\dot{A}^{G_{0}}\in\operatorname{HS}^{G_{0}}=\mathcal{N}$. ∎
@@
Suppose toward a contradiction that there exist an ordinal $\theta$, a name $\dot{f}\in\operatorname{HS}$, and a condition $p_{0}\in\mathbb{P}$ such that
@@
Remark 3.22 (Countable subsets of $A$).
In contrast with the finite-support generalized Cohen model (where the analogous $A$ can be Dedekind-finite; cf. [16, §2.4]), the present countable-support filter admits many canonical countable subsets of $A$. Fix $E\in[\omega_{1}]^{\leq\omega}\cap V$ and set
$\dot{A}_{E}\;:=\;\{\langle\dot{c}_{\alpha},\mathbf{1}_{\mathbb{P}}\rangle:\alpha\in E\}.$
Then $\operatorname{Fix}(E)\leq\operatorname{Stab}(\dot{A}_{E})$, hence $\dot{A}_{E}\in\operatorname{HS}$ and
$A_{E}:=\dot{A}_{E}^{G_{0}}=\{c_{\alpha}:\alpha\in E\}\in\mathcal{N}.$
If $E$ is infinite and $b:\omega\to E$ is a bijection in $V$, then the sequence $\langle c_{b(n)}:n\in\omega\rangle\in\mathcal{N}$ (Lemma 3.23) witnesses that $A$ is Dedekind-infinite in $\mathcal{N}$.
@@
Suppose $h$ is an order automorphism of $(X,\prec)$ and $h\neq\mathrm{id}_{X}$. Let $x_{0}$ be the $\prec$-least element of $X$ with $h(x_{0})\neq x_{0}$.
If $h(x_{0})\prec x_{0}$, then by minimality of $x_{0}$ we have $h(h(x_{0}))=h(x_{0})$. But order preservation gives $h(h(x_{0}))\prec h(x_{0})$, contradiction.
If $x_{0}\prec h(x_{0})$, apply the previous case to $h^{-1}$, which is also an order automorphism of $(X,\prec)$, and satisfies $h^{-1}(x_{0})\prec x_{0}$, again a contradiction.
Therefore $h=\mathrm{id}_{X}$. ∎